OpenStudy (anonymous):

The sum of n terms of a series is given by the expression, \[S_n = 3(-2)^n+5\] Determine the 3rd term in the series. ------------------------------- If I'm correct, the first term can be found by inserting n=1 as the sum of the first term. That gives me -1. Then I found the sum of two terms using the above formula, giving me -1. Then, \[a+ar=-1\] \[(-1)+(-1)(\frac{-9}{8})\] to find r. Yet it doesn't give me the right answer, as I found three individual terms, added them up, and they give a different number than the given formula.

OpenStudy (amistre64):

the third partial sum? or the third term of the sequence?

OpenStudy (amistre64):

and does n start at 0 or 1? different texts have different ideas for that as well

OpenStudy (anonymous):

Third term of the sequence. As for n, it does not specify. I tried both n=0 and n=1 to no avail.

OpenStudy (amistre64):

im wondering then if its talking about the initial terms that are being summed up

OpenStudy (amistre64):

s1 = a1 by default for the first term

OpenStudy (amistre64):

s2 = s1 + a2 a2 = s2 - s1

OpenStudy (amistre64):

s3 = s1 + s2 - s1 + a3

OpenStudy (amistre64):

a3 = s3 - s2 is what it might turn out to be

OpenStudy (amistre64):

does that make sense?

OpenStudy (anonymous):

Going by your last equation, I believe you are subtracting adjacent sums so it leaves one 'left over' which would be the nth term. I believe I have tried this before, but I will try it again.

OpenStudy (amistre64):

since youve tried the partial sum approach to no avail, this is the only thing that would make sense for me

OpenStudy (anonymous):

Hmm, tried both n=0 and n=1 for the first term and then used ar^2 or ar = the number I got that is supposedly the term to find a, and then working forwards, my sums are still different.

OpenStudy (amistre64):

i get -36 for an answer a1 = -1 a2 = 18 a3 = -36 --------- s3 = -19 but s3 = 3(-2)^3 + 5 = -21 hmmm

OpenStudy (anonymous):

Same here... I tried a=8 (n=0) as well to no avail.

OpenStudy (amistre64):

-1 + a2 + a3 = -21 , assuming n=1,2,3,... s2 = -1 + a2 = 3(-2)^2+5 = 3(4)+5 = 17 -1 + a2 = 17, a2 = 18 -1+18 + a3 = -21 a3 = -21+1-18 = -38

OpenStudy (amistre64):

so my mathing was off the first time, but the idea was sound enough lol

OpenStudy (amistre64):

you might wanna check that for n starting at 0 just in case tho

OpenStudy (amistre64):

in which case a3 = a2 = 18

OpenStudy (anonymous):

Yes, -38 seems to work. I'm still trying to figure out why my original method of ar = 17 didn't work though.

OpenStudy (amistre64):

probably due to some inherent flaw in the concept ;)

OpenStudy (anonymous):

Do you know the flaw? I kinda think I need to use two equations to solve it, since one doesn't really prove anything except for that case.

OpenStudy (amistre64):

you say the sum of 1 term is -1 then you say the sum of the first 2 terms is ... -1 again for some reason

OpenStudy (amistre64):

and it looks like your trying to fit this into a geometric sequence, which itself could be a mistake to begin with

OpenStudy (anonymous):

Oops I meant a + ar=17. I just assumed it was a geometric sequence because that's what I'm studying right now... Anyways, thanks for showing me your method. It makes sense and it works, so I'm no complaining :)

OpenStudy (amistre64):

youre welcome, and good luck :)

OpenStudy (anonymous):

Wait, doesn't S3 = 3(-2)^3+5 = -19, not -21?

OpenStudy (kinggeorge):

\(S_3=3(-2)^3+5=3(-8)+5=-24+5=-19\) So @amistre64's first attempt was correct I think. I am also getting \(a_3=-36\).

OpenStudy (anonymous):

@KingGeorge - Which approach did you use? Subtracting S3-S2?

OpenStudy (kinggeorge):

\(S_1=-1\), so \(a_1=-1\) \(S_2=17\) so \(a_2=S_2-a_1=17-(-1)=18\) \(S_3=-19\) so \(a_3=S_3-(a_1+a_2)=-19 -(18-1)=-19-17=-36\)

OpenStudy (kinggeorge):

In short, \(S_3-S_2\)

OpenStudy (anonymous):

Okay, thank you!

OpenStudy (kinggeorge):

You're welcome.